2sinxcosx+sinx-2cosx-1=0

solve in interval [0,2)
1
Expert's answer
2012-01-26T11:36:31-0500
цу understand that correct problem is solve in interval [0,2pi) but if realy
solve in interval [0,2) then you can choose from results that values that need(
but for [0;2) you get no answer it will be empty set)

we can rewrite it as
2cos(x)(sinx-1)+(sinx-1)=0 or (2cosx+1)(sinx-1)=0
so cosx=-1/2 or sin x=1 and
we have x=5pi/6 or x=7pi/6 or x=pi/2